OpenStudy (mathmagician):

Can an irrational power of an irrational number be rational? (complex numbers should be omitted)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

proof consider \[\sqrt{2}^{\sqrt{2}}\] if it is rational, then you know of one if it is irrational then consider \[(\sqrt{2}^{\sqrt{2}})^{\sqrt{2}}=\sqrt{2}^2=2\]which is rational in either case there must be one